Prev Next

SW&IT Meeting Minutes 2021-07-19

Attendees: David, Matthieu, Michael, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • xxx
  • This week
    • xxx

YZ

  • Previous:
    • Some work on TMS/OSCO.
    • Worked with Paul on TCS/OSS and Swingarm tests.
  • Next:
    • Continue to work on some aspects of TMS/OSCO.
    • Continue to work with Paul on OSS and Swingarm tests and patches.

PSG

  • Last week
    • OSS was patched on tcs1 to run the swings, both ARGOS and standard.
    • D425 unit could not be reached, it has reverted back to its old settings.
    • This behavior was tested in the lab, and the settings were preserved over a power cycle, so this is a mystery.
    • Was able to test the existing OSS functionalty.
    • Fixed some bugs that were introduced by adding ARGOS swing arms.
  • This week
    • Gustavo will try to update the D425 when he is back from vacation.
    • Testing white light source from OSS, since it is reachable for tcs1.
Topic revision: r2 - 26 Jul 2021, PaulGrenz
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback